Solve the following system of equations
x1 + x2 + x3 = 6
3x1 – 4x2 – 2x3 = 1
3x1 – 2x2 + x3 = 2

Answer by Alan3354(69443)

x1 = 47/11
x2 = 46/11
x3 = -27/11
--------
I use determinants to solve these.
I made an Excel sheet that does the solution by use of determinants.